### v2

# Learning to Rank Critical Road Segments via Heterogeneous Graphs with OD Flow Integration

Abstract
Existing learning-to-rank methods for road networks often fail to incorporate origin–destination (OD) flows and route information, limiting their ability to model long-range spatial dependencies. To address this gap, we propose HetGL2R, a heterogeneous graph learning framework for ranking road-segment importance. HetGL2R builds a tripartite graph that unifies OD flows, routes, and network topology, and further introduces attribute-guided graphs that elevate node attributes into explicit nodes to model functional similarity. A heterogeneous joint random walk algorithm (HetGWalk) samples both graph types to generate context-rich node sequences. These sequences are encoded with a Transformer to learn embeddings that capture long-range structural dependencies driven by OD demand and route configuration, as well as functional associations derived from attribute similarity. Finally, a listwise ranking strategy with a KL-divergence loss evaluates and ranks segment importance. Experiments on three SUMO-generated simulated networks of different scales show that, against state-of-the-art methods, HetGL2R achieves average improvements of approximately 7.52%, 4.40% and 3.57% in ranking performance.
keywords:
Learning to Rank , Heterogeneous Graph , Random Walk , Ranking , Road Networks
[…]
[th]organization=School of Software, Liaoning Technical University, postcode=125105, city=Huludao, country=China
[…]
[pku]organization=Faculty of Electronic and Information Engineering, Liaoning Technical University, postcode=125105, city=Huludao, country=China
 [^1] Our code and data are available at : https://github.com/iCityLab/HetGL2R.
1 Introduction
Efficient and resilient road networks are essential for ensuring smooth urban mobility and public safety. When a single road segment becomes congested or blocked, the resulting disruption often propagates along multiple routes, leading to large-scale delays or even citywide paralysis. Therefore, identifying critical road segments—those whose failure would significantly degrade overall network performance—is of great importance for traffic management and infrastructure planning (rf1).
Figure 1: A small simulated road network. The network contains 11 nodes and two OD pairs: $v_{1}$, $v_{4}>$ with 300 vehicles on one path, and $v_{8}$, $v_{7}>$ with 500 vehicles on two paths.
Prior studies on road-network criticality were primarily based on complex network theory, in which centrality measures (e.g., early betweenness centrality (betweenness) and PageRank (rf3), and more recent methods such as IDME (IDME), EPC (EPC), OVED-Rank (rf44), LSS (rf42), and GLC (rf41)) were used to quantify the structural importance of nodes. These approaches are intuitive and easy to interpret but fail to incorporate the rich attribute features and dynamic traffic behaviors associated with each road segment. In reality, a segment’s criticality depends on multiple factors such as traffic volume, number of lanes, and functional hierarchy, all of which are neglected in purely topological metrics. Some studies have attempted to combine centrality measures with attribute features (e.g., study (rf3)), but most rely on manually tuned parameters, limiting their scalability and generalization to large real-world networks.
In recent years, graph neural network (GNN)–based learning methods (rf4; rf8) have attracted increasing attention in the field of node importance assessment. These models integrate node attributes with local topological information through message-passing mechanisms to learn effective node embeddings. However, due to the limited receptive field of GNNs, they struggle to capture the long-range path dependencies commonly present in traffic flows, which limits their performance in critical road segment ranking tasks.
In transportation networks, the global influence of a road segment is predominantly governed by origin–destination (OD) flows, which quantify the travel demand between each OD pair. The OD flows determine which paths are utilized and how traffic load is distributed, thereby shaping each segment’s functional role and criticality (od1; od2; od3). For instance, if a segment lies on the shortest paths of multiple high-demand OD pairs, any incident or restriction there can induce congestion that propagates upstream along the affected OD paths, leading to severe network-wide delays. Conversely, a peripheral segment carrying a similar traffic volume but serving only a few OD flows would have a far more limited systemic impact if disrupted.
As illustrated in Figure 1, a simulated road network contains two OD flows, $v_{1}$, $v_{4}>$ and $v_{8}$, $v_{7}>$. When an incident occurs at node $v_{11}$, congestion propagates upstream along the path of OD flow $v_{8}$, $v_{7}>$ to nodes $v_{10}$ and $v_{9}$. This forces some travelers to reroute, thereby increasing the congestion risk on distant segments $v_{5}$ and $v_{6}$. In contrast, segments $v_{2}$ and $v_{3}$ remain largely unaffected as they do not carry the affected OD flows. This example highlights that the influence among road segments is fundamentally governed by the OD flows they serve and the paths they form. Conceptually, spatial dependencies in a road network stem from functional interactions among OD flows, paths, and segments, rather than from physical connectivity. Consequently, the OD-dominated functional relationships are pivotal for evaluating node criticality.
Existing graph learning approaches face two fundamental limitations in this task. First, they primarily learn structural representations from the physical topology of road networks while overlooking the functional structures shaped by OD flows. Second, their local neighborhood aggregation mechanism is constrained by network depth, resulting in limited receptive fields. As a result of these limitations, they struggle to capture long-range spatial dependencies and consequently fail to accurately model the impact patterns of local disruptions (e.g., an incident at node $v_{11}$)—where the scope of impact is dictated by the functional roles of segments within OD flows and paths, rather than by physical proximity. Although CRRank (rf1) incorporates OD information to evaluate the functional importance of road segments, its reliance on manually designed aggregation rules and similarity metrics constrains its expressiveness and effectiveness, as it cannot learn complex interaction patterns in an end-to-end manner.
To overcome these limitations, this paper proposes HetGL2R, a heterogeneous graph learning framework that integrates OD flow information for road-segment ranking. The key contributions of this study can be summarized as follows:
  • 1.
    Methodological Innovation: HetGL2R constructs a tripartite graph (connecting OD pairs, paths, and road segments) to integrate the functional structure of the road network with its physical structure. Simultaneously, it introduces three attribute-node bipartite graphs—one each for OD pairs, paths, and segments—to explicitly model the attribute-driven functional similarities. A joint random walk mechanism operates on these graphs collectively to sample context-rich node sequences, which are then processed by a Transformer encoder to learn the long-range dependencies and functional roles of segments for accurate criticality ranking.
  • 2.
    Theoretical Insight: Our core insight is that spatial dependencies in traffic networks are governed by functional structure rather than physical topology. To capture this structure, we introduce a joint random walk that integrates two complementary perspectives: functional role relationships (OD-path-segment) and attribute-driven functional similarities. Consequently, the model learns the underlying mechanisms of spatial dependency, thereby enabling accurate assessment of nodes’ spatial influence.
  • 3.
    Practical Implication: Extensive evaluation on three simulated traffic network datasets of varying scales demonstrates that our proposed framework consistently outperforms existing baseline methods. The framework can serve as a foundation for applications such as transportation resilience analysis, emergency route planning, and infrastructure maintenance prioritization.
2 Related Works
In this section, we discuss three research areas that are closely related to this paper.
2.1 Evaluation of node importance in networks
Existing research on identifying critical nodes in networks can be broadly categorized into two approaches: centrality-based methods and machine learning-based methods. The limitation of centrality-based methods lies in their inability to fully leverage the multidimensional features of nodes. Although study (rf3) attempts to improve critical node identification in road networks by integrating node centrality with attribute features, its performance heavily depends on manually set topology-feature trade-off parameters. This leads to a lack of generalizability in quantifying the contributions of topology and features, limiting its applicability in large-scale real-world road networks. To address these limitations, LSS (rf42) estimates node importance based on local information, eliminating the need for global knowledge and complex parameter tuning. Study (rf41) proposed a dual-perspective approach that identifies influential nodes by considering inter-cluster connectivity and local neighborhood importance. However, similar to traditional centrality-based methods, these approaches still fail to incorporate node attribute features or dynamic network behaviors, limiting their applicability in more complex real-world scenarios such as road networks. In recent years, researchers have gradually shifted towards machine learning-based methods, which can be divided into two main technical approaches: reinforcement learning frameworks (e.g., FINDER proposed in study (rf4), which utilizes GraphSAGE (rf5) to learn node importance; study (rf6), which combines attribute-aware graph attention networks with double DQN to optimize critical node identification in sparse graphs) and graph representation learning (e.g., TraNode2vec from study (rf7), which assesses node importance by integrating topological embedding clustering and traffic flow; study (rf8), which introduces the MGL2Rank framework that employs multi-graph fusion techniques to jointly model topology and road segment attributes). However, when applying these methods to road network node importance evaluation, they fail to incorporate OD flow and path selection. As we analyzed in the introduction, these two types of information are crucial for assessing node importance in road networks. Therefore, there remains significant room for improvement in the performance of these methods.
2.2 Learning to rank
Learning to rank (LTR) originated in the field of information retrieval and is a technique for ordering documents, items, or other entities based on their relevance. In recent years, it has been widely applied in various domains, including recommender systems (rf9), sentiment analysis (rf10), and object detection (rf11). Depending on the characteristics of the input and output spaces, LTR methods can be broadly categorized into three types: pointwise, pairwise, and listwise approaches. The pointwise approach (rf12) treats each item independently, formulating the ranking task as a classification or regression problem. However, this method neglects the relative relationships among items. In contrast, the pairwise approach (rf13) transforms the ranking task into a binary classification problem by comparing item pairs to determine their relative order. Prior research (rf14) has introduced graph neural networks into pairwise ranking to approximate the importance of nodes within a network. Although the pairwise method offers improved accuracy over the pointwise method, it still fails to fully capture the global context and relative importance of items within the entire set, limiting its overall performance. The listwise approach (rf15) treats ranking as a holistic optimization problem and has become a focal point of research due to its modeling and computational complexity. For example, one study (rf16) analyzes how list length and loss function selection affect algorithm stability through generalization error bounds, while another (rf17) creatively reformulates ranking as a multi-task regression problem, overcoming the limitations of traditional linear models. With the rise of deep learning, researchers (rf18; rf19; rf20) have incorporated attention mechanisms and recurrent neural networks into listwise modeling, significantly enhancing performance in scenarios involving complex interactions. On the application side, DiAL (rf21) is the first to integrate the $\alpha$-NDCG diversity metric into listwise loss functions. Further advancements include leveraging large language models to generate reference answer sequences for listwise evaluation (rf22; rf45), as well as developing differentiable approximations to ranking indicators that closely match true loss functions with limited data (rf23). Recent work (rf24) also proposes a regression alignment strategy to address the multi-objective trade-off between score calibration and ranking optimization. Building upon the core principles of listwise ranking, this paper proposes a heterogeneous graph learning framework tailored for node ranking in transportation networks.
2.3 Heterogeneous graph learning
Heterogeneous graphs (rf25) model multiple types of nodes and edges, integrating information from heterogeneous data sources, thereby demonstrating a stronger representation ability for complex systems. Representation learning on heterogeneous graphs aims to obtain meaningful node embeddings to support various downstream tasks, such as node ranking and link prediction (rf26; rf27). Many studies have focused on designing graph neural network models suitable for heterogeneous graphs. For example, most of these networks (rf28; rf29; rf30) utilize meta-paths and type-aware aggregation strategies to capture the heterogeneity of the graph effectively, that is, the difference in the type of nodes and edges. HetGNN (rf31) combines random walk strategies with bidirectional LSTM structures, effectively integrating heterogeneous node attributes and their neighboring information. HetSANN (rf32) directly encodes the information of different types of nodes and edges through a type-aware attention mechanism, enhancing the model’s ability to represent the heterogeneous graph structure. These methods are distinctive in their modeling strategies, but most still rely on message-passing mechanisms, which do not fully exploit the potential structural relationships between nodes in heterogeneous graphs. To address this, M2CHGNN (rf43) employs a dual-view encoding strategy that integrates attention mechanisms with contrastive learning to generate high-quality node representations for heterogeneous graphs. UI-HEHo (rf47) effectively learns high-order node embeddings via multi-head attention mechanisms and multi-task shared features. SchemaWalk (rf34) proposes an unbiased edge-type exploration strategy, providing finer-grained and more generalizable methods for embedding heterogeneous networks. JUST (rf35) introduces the ”Jump-Stay” strategy for adaptive node sampling, further improving the representation capability. HMSG (rf36) constructs a meta-path-based multi-relational graph neural network, and employs contrastive learning together with Deep Graph Infomax (DGI) (rf37) to jointly learn high-quality node representations. HPSRW (rf38) combines higher-order Markov chains and the Spacey Walk method to effectively mitigate the stability issues in meta-path sampling. While heterogeneous graph learning has advanced in various domains, existing general-purpose models lack specialized designs for the unique characteristics of transportation networks. Even when applied to our constructed graph that explicitly incorporates OD flows and route information, these models still fail to effectively leverage its structural properties, leading to limited capability in capturing the spatial dependencies inherent in road networks. To the best of our knowledge, CRRank (rf1) is the first work to apply heterogeneous graphs to transportation network analysis, achieving node ranking by integrating OD flow and path information. However, this method relies on manually defined computational rules and lacks end-to-end learnability, leaving room for substantial improvements in model performance.
3 Preliminary
3.1 Definitions and problem statement
Before presenting our method, we first provide the definitions relevant to this work.
Definition 1 (Road Network): A road network is a directed graph $G(V^{l},{\color[rgb]{0,0,0}E^{l}})$, where $V^{l}$ is the set of nodes; $E^{l}$ is the set of directed edges; An edge $e_{i,j}$ indicates that node $v^{l}_{i}$ is the incoming neighbor of node $v^{l}_{j}$. Each node $v^{l}_{i}$ has an initial attribute vector. In this paper, $v^{l}_{i}$ specifically refers to a road segment or a grid.
Definition 2 (OD Flow): An OD flow is represented as a triple $\textlangle v_{c},v_{t},f\textrangle$, where $v_{c}\in V^{l}$ denotes the origin node, $v_{t}\in V^{l}$ denotes the destination node, and $f$ represents the number of vehicles traveling from $v_{c}$ to $v_{t}$ within a specified time period. For instance, the road network in Figure 1 contains two OD flows: $v_{1},v_{4},300>$ and $v_{8},v_{7},500>$.
Definition 3 (path): A path $p:\textlangle v^{l}_{1},v^{l}_{2},...,v^{l}_{|p|}\textrangle$ is a sequence of nodes in the road network, encompassing all the nodes traversed during a trip associated with a specific OD flow. Here, $|p|$ denotes the number of nodes in $p$. Evidently, an OD flow can be associated with multiple paths. For example, the OD flow $v_{8},v_{7}>$ corresponds to two paths: $v_{8},v_{9},v_{5},v_{6},v_{7}>$ and $v_{8},v_{9},v_{10},v_{11},v_{7}>$.
Definition 4 (Trip Graph): A trip graph is a heterogeneous graph $TG(V,A,M)$, where $V=V^{o}\cup V^{p}\cup V^{l}$ consists of OD nodes $(V^{o})$, path nodes $(V^{p})$, and road segment nodes $(V^{l})$ with corresponding attribute matrices $A^{o}$, $A^{p}$, and $A^{l}$ encoding features. The adjacency structure includes (1) $M_{o,p}$ representing traffic proportions between OD pairs and paths $(m_{o,p}\in[0,1])$, (2) $M_{p,l}$ as a binary path-segment incidence matrix $(m_{p,l}=\mathbb{I}(v^{l}\in p))$ and (3) a topologically weighted segment connectivity matrix $M_{l,l}$ defined as
$m_{i,j}=\left\{\begin{array}[]{c}\frac{\sum_{p\in\mathcal{P}_{i,j}}e^{-d_{i,j}^{p}}}{\sum_{p\in\mathcal{P}_{i}}\sum_{v_{k}^{l}\in p}e^{-d_{i,k}^{p}}}\text{ if }\exists p\text{ where }v_{i}^{l}\prec_{p}v_{j}^{l}\\ 0\quad\text{ otherwise }\end{array}\right$. (1)
where $\mathcal{P}_{i,j}=\left\{p\in V^{p}\mid v_{i}^{l},v_{j}^{l}\in p,v_{i}^{l}\prec_{p}v_{j}^{l}\right\}$ denotes paths containing both segments with $v_{i}^{l}$ upstream of $v_{j}^{l}$, $d_{i,j}^{p}$ counts intermediate segments between them in path $p$, $e$ is the propagation factor, and $\mathcal{P}_{i}=\left\{p\in V^{p}\mid v_{i}^{l}\in p\right\}$ is the set of paths containing $v_{i}^{l}$. Finally, the adjacency matrix $M_{p,l}$ and attribute matrices $A^{o}$, $A^{p}$, and $A^{l}$ are row-normalized using the L1 norm, resulting in $\overline{M}_{p,l}$, $\overline{A}^{o}$, $\overline{A}^{p}$, and $\overline{A}^{l}$, respectively.
While the Trip Graph (TG) effectively integrates the macro-level functional organization and topological structure of the road network, it struggles to capture local, attribute-driven semantic similarities among nodes of the same type. For instance, the functional similarity between OD pairs that both connect residential and commercial areas, or between road segments sharing comparable attributes like the number of lanes or traffic capacity, is crucial for learning comprehensive and robust segment representations. To address this limitation, we introduce the Attribute-guided Graph (AG) as a complementary structure to the TG at the semantic level. Its core innovation lies in elevating node attributes to the status of explicit nodes, thereby constructing a heterogeneous bipartite graph that directly models these attribute-driven functional relationships.
Definition 5 (Attribute-Guided Graph): For each node type $\phi\in\left\{o,p,l\right\}$, an attribute-guided graph is defined as a weighted bipartite graph $AG_{\phi}(V^{\phi}\cup U^{\phi},A^{\phi})$, where $V^{\phi}$ is the set of nodes of type $\phi$, $U^{\phi}$ is the set of attribute nodes associated with type $\phi$, $A^{\phi}$ is the set of weighted edges between nodes $v_{i}^{\phi}\in V^{\phi}$ and attribute nodes $u_{k}^{\phi}\in U^{\phi}$. The edge weight between $v_{i}^{\phi}$ and $u_{k}^{\phi}$ is denoted as $a_{i}^{\phi}\in A^{\phi}$. Special cases: For OD nodes $(\phi=0)$, $AG_{o}(V^{o}\cup U^{o},A^{o})$. For paths $(\phi=p)$, $AG_{p}(V^{p}\cup U^{p},A^{p})$. For road segments $(\phi=l)$, $AG_{l}(V^{l}\cup U^{l},A^{l})$.
[…]
Table 1: Notation and their definitions.
Symbol Meaning
$G=(V^{l},E^{l})$ Road network, where $E^{l}$ represents the set of directed edges.
$V^{o},\,V^{p},\,V^{l}$ Sets of OD nodes, paths, and road segments, respectively.
$A^{o},\,A^{p},\,A^{l}$ Attribute matrices for OD nodes, paths, and road segments, respectively.
$U^{o},\,U^{p},\,U^{l}$ Sets of attribute nodes for OD nodes, paths, and road segments, respectively.
$\langle v_{c},v_{t},f\rangle$ OD flow: $f$ vehicles traveling from origin $v_{c}$ to destination $v_{t}$ within a time period.
$p:\langle v^{l}_{1},v^{l}_{2},\dots,v^{l}_{|p|}\rangle$ A path composed of $|p|$ road segments $v^{l}_{1},v^{l}_{2},\dots,v^{l}_{|p|}$.
$TG=(V,A,M)$ Trip Graph, where $V=V^{o}\cup V^{p}\cup V^{l}$, $A=A^{o}\cup A^{p}\cup A^{l}$, and $M=M_{o,p}\cup M_{p,l}\cup M_{l,l}$.
$M_{o,p}$ OD–Path traffic proportion matrix, with entries $m_{o,p}\in[0,1]$.
$M_{p,l}$ Path–Segment incidence matrix, where $m_{p,l}=\mathbb{I}\!\left(v^{l}\in p\right)$.
$M_{l,l}$ Topologically weighted segment–segment connectivity matrix.
$AG_{\phi}\!\left(V^{\phi}\!\cup\!U^{\phi},\,A^{\phi}\right)$ An attribute-guided graph for a specific node type $\phi$, where $\phi\in\{o,\,p,\,l\}$ corresponds to OD nodes, paths, and road segments, respectively.
$P_{i,j}^{TG}$ The transition probability from node $v^{\phi}_{i}$ to node $v^{\phi}_{j}$ on the TG, where $v^{\phi}_{i},\,v^{\phi}_{j}\in V$, and $\phi\in\{o,\,p,\,l\}$.
$P^{AG_{\phi}}_{i,k}$ The transition probability from node $v^{\phi}_{i}\in V^{\phi}$ to attribute $v^{\phi}_{k}\in U^{\phi}$ on the $AG_{\phi}$, where $\phi\in\{o,\,p,\,l\}$.
$Y=[y_{1},y_{2},\dots,y_{k}]$ Ground-truth node ranking list, where $k$ is the list length.
$\hat{Y}=[\hat{y}_{1},\hat{y}_{2},\dots,\hat{y}_{k}]$ Predicted node ranking list, where $k$ is the list length.
$\alpha\in[0,1]$ Walk bias factor in the TG.
$\beta\in[0,1]$ Probability of selecting a depth-first exploration strategy at each step of the random walk.
numnum, lenlen Parameters for generating numnum random walk sequences, each of length lenlen.
$h_{i}$ Final representation vector for a road segment.
Table 1 provides a summary of key notation and their definitions. This paper proposes a joint random walk-based representation learning framework on the TG and AG to improve the performance of the road segment ranking. This is achieved by integrating OD-driven functional structure and attribute-based functional similarity into the learned representations, which enables our method to more accurately capture long-range spatial dependencies in the traffic network.
3.2 Research objectives
(RO1) To design a joint heterogeneous graph random walk mechanism that generates context-rich sequences by sampling from both OD-driven functional structures and attribute-based similarities.
(RO2) To formulate a learning-to-rank framework that leverages the functional structure of road networks to achieve accurate prioritization of the most critical segments.
(RO3) To investigate whether the proposed model’s performance on road-network ranking tasks—relative to topology-based baselines—stems from its capacity to capture the underlying functional structure.
4 Methodology
This paper proposes HetGL2R, a deep learning framework (Figure 2) based on random walks on heterogeneous graphs, to improve traffic network ranking by integrating their functional structures with physical topology. The framework first extracts information from vehicle trajectory data to construct an OD-path-segment tripartite graph, which integrates the functional and topological structures of the road network into a unified representation. Simultaneously, a set of three attribute-guided bipartite graphs—one each for OD pairs, paths, and segments—are constructed to model functional similarities among nodes. Subsequently, the heterogeneous graph joint random walk algorithm, HetGWalk, is applied to these graphs to sample context-rich node sequences. These sequences are encoded using a Transformer encoder to learn road segment embeddings that capture long-range dependencies. Finally, a ranking learning module comprehensively evaluates and ranks the importance of road segments.
Figure 2: Framework of the proposed method.
4.1 HetGWalk algorithm
4.1.1 HetGWalk Overview
This subsection introduces the joint random walk algorithm, HetGWalk. The algorithm generates node sequences through probabilistic sampling on both the Trip Graph (TG) and the Attribute-guided Graph (AG): walks on the TG utilize OD-flow-path-segment connections to capture functional relationships and local topology of traffic networks, while walks on the AG leverage node-attribute relationships to encode functional similarities. The AG is introduced to enrich node context based on attributes, which increases sequence diversity and reduces the walks’ tendency to over-sample high-centrality nodes. By integrating these two types of walks, HetGWalk produces heterogeneous sequences that integrate functional and topological information, providing rich context for downstream representation learning. Figure 3 illustrates the sampling process.
The walk starts from the OD node and traverses all heterogeneous nodes in TG and all three AGs. To simultaneously capture both the local and global structures of TG, HetGWalk integrates depth-first exploration and breadth-first exploration. Given the current node $v^{\phi}_{i}$, we flip a coin with probability $\alpha$ of landing heads. If it lands on heads, we step in TG, with the probability of transitioning to $v^{\phi}_{j}$ given by:
$P_{i,j}^{TG}=\beta\tilde{P}_{i,j}^{TG}+(1-\beta)\tilde{P}_{i^{-},j}^{TG}$ (2)
where $\phi\in\{o,\,p,\,l\}$ denotes the node type, corresponding to OD nodes, paths, and road segments, respectively; $\beta$ is the probability of choosing depth-first exploration at each step of the walk; $i^{-}$ is the index of the predecessor node of $v^{\phi}_{i}$; $\tilde{P}_{i,j}$ represents the transition probability from $v^{\phi}_{i}\in V$ to $v^{\phi}_{j}\in V$ in a depth-first manner, which is calculated as follows:
$\tilde{P}_{i,j}^{TG}=\left[\begin{array}[]{ccc}0&M_{op}&0\\ 0&0&\bar{M}_{pl}\\ 0&(1-\varepsilon)\bar{M}_{pl}^{T}&\varepsilon M_{ll}\end{array}\right]\in\mathbb{R}^{|V|\times|V|}$ (3)
where, $\varepsilon$ denotes the probability of selecting segment nodes for the next step from the current segment node. If the coin lands tails up, the walk proceeds on the AG corresponding to the current node type in two steps:
The walk first transitions from the current entity node $v^{\phi}_{i}$ to one of its associated attribute nodes $u^{\phi}_{k}$. The probability of this transition, $P_{i,k}^{AG_{\phi}}$, is proportional to the importance of attribute $u^{\phi}_{k}$ to node $v^{\phi}_{i}$, specifically given by the normalized attribute value $\bar{a}^{\phi}_{k,i}$:
Figure 3: Joint random walk sampling procedure.
$P_{i,k}^{AG_{\phi}}=\bar{a}^{\phi}_{k,i}$ (4)
This means that attributes contributing more to node $v^{\phi}_{i}$ have a higher probability of being sampled.
Next, the walk transitions from this attribute node $u^{\phi}_{k}$ to a new entity node $v^{\phi}_{j}$. The probability of this transition, $P_{k,j}^{AG_{\phi}}$, is determined by the attribute-value similarity between the target node $v^{\phi}_{j}$ and the previous entity node $v^{\phi}_{i}$ regarding attribute $u^{\phi}_{k}$. We employ a similarity measure based on the absolute difference:
$P_{k,j}^{AG_{\phi}}=sim_{i,j}^{k}=1-\left|\bar{a}^{\phi}_{k,j}-\bar{a}^{\phi}_{k,i}\right|$ (5)
1
Input: TG, AG, $\alpha$, $f$, num, len, $\beta$
Output: The sampling sequence collections of $|V^{o}|$ nodes
2
Data: L1 norm to row-normalize $M^{T}$ and $A^{T}$ /* walk starts from an OD node */
3
4foreach $v^{o}_{t}\in V^{o}$ do
5 $S_{t}=\left\{\right\}$;
6 for $n=1$ to num do
7    for $l=1$ to len do
8       $s=\{v^{(t)}\}$;
9       $i\leftarrow t$;
10         flip a coin with probability $\alpha$ of landing heads;
11       if heads then
12            With probability $P_{i,j}^{TG}$, move to node $v^{(j)}$ in TG;
13            Append $v^{(j)}$ to $s$;
14
15       else
16            With probability $P^{AG_{\phi}}_{i,k}$, move to $u^{\phi}_{k}\in U^{\phi}$ in $AG_{\phi}$;
17            Append $u^{\phi}_{k}$ to $s$;
18            With probability $P^{AG_{\phi}}_{k,j}$, move to $v^{(j)}\in V^{\phi}$ in $AG_{\phi}$;
19            Append $v^{(j)}$ to $s$;
20
21       $i\leftarrow j$;
22
23     Append $s$ to $S_{t}$;
24
25
26return $\{S_{1},S_{2},\dots,S_{|V^{o}|}\}$;
Algorithm 1 HetGWalk
This design encourages the walk to jump to nodes that are similar to the origin node $v^{\phi}_{i}$ in terms of the specific attribute $u^{\phi}_{k}$, thereby capturing functional similarity even if the two nodes are not directly connected in the original Trip Graph.
Through this mechanism, the attribute-guided walk incorporates both the influence and similarity of node attributes into the transition process. The overall probability of reaching $v^{\phi}_{j}$ from $v^{\phi}_{i}$ is computed as a weighted aggregation over all possible paths passing through attribute nodes $u^{\phi}_{k}$, effectively constructing an implicit similarity network enriched with attribute semantics. The implementation details of HetGWalk are described in Algorithm 1.
Line 1: The algorithm iterates over each OD node $v^{o}_{t}$ in the set $V^{o}$, which serves as the starting point for random walks. Line 2: Initializes an empty collection $S_{t}$ to store the generated sequences for the current OD node $v^{o}_{t}$. Line 3: For each starting node, the algorithm generates numnum independent random walk sequences. Line 4: Each sequence grows incrementally until reaching the predefined length lenlen. Line 7: At each step, a random decision is made: with probability $\alpha$, the walk continues in the Trip Graph; otherwise, it proceeds to the Attribute-guided Graph. Lines 9-10: If the decision selects TG traversal, the walk moves to a neighboring node $v^{(j)}$ according to the transition probability $P_{i,j}^{TG}$ (Eq. 2) and appends it to the sequence.
Lines 12-15: If AG traversal is selected, the walk executes a two-stage process: First transitions from the current node $v^{(t)}$ to an attribute node $u^{\phi}_{k}$ using probability $P_{i,k}^{AG_{\phi}}$ (Eq. 4). Then immediately moves to entity node $v^{(j)}$ using probability $P_{k,j}^{AG_{\phi}}$ (Eq. 5). Both the attribute node and the new entity node are appended to the sequence. Line 16 Updates the current node position to continue the walk from $v^{(j)}$. Line 17: Stores the completed sequence $s$ in the collection $S_{t}$. Line 18: Returns all generated sequences for all starting OD nodes.
4.1.2 Theoretical Properties of HetGWalk
To rigorously analyze the behavior of our proposed heterogeneous graph representation learning framework, we formalize the properties of attribute-guided graphs (AGs) and joint random walks on the trip graph (TG) and AGs.
Definition 7 (Type-Specific Attribute Propagation Operators): For a node $v_{i}^{\phi}\in V^{\phi}$, $\phi\in\left\{o,p,l\right\}$, the propagation operator $\rho_{i}^{\phi}(j)$ measures the probability of influence propagation from $v_{i}^{\phi}$ to another node $v_{j}^{\phi}$ via attributes:
$\color[rgb]{0,0,0}\rho_{i}^{\phi}(j)=\sum_{k\in U^{o}}\bar{a}_{k,i}^{\phi}\cdot\left(1-\left|\bar{a}_{k,j}^{\phi}-\bar{a}_{k,i}^{\phi}\right|\right),\quad v_{j}^{\phi}\in V^{\phi}$} (6)
This operator captures type-specific semantic dependencies within the attribute-guided graph.
Theorem 1 (Type-Aware Dynamic Semantic Propagation): For any node type $\phi\in\left\{o,p,l\right\}$, attribute-guided random walk transitions on $AG_{\phi}$ implement a type-specific semantic propagation process. Specifically, when the walk is currently at node $v_{i}^{\phi}$, the total transition probability from $v_{i}^{\phi}$ to any node $v_{j}^{\phi}$ is exactly equal to the propagation operator $\rho_{i}^{\phi}$.
Proof: By the AG walk rules, the probability of moving from $v_{i}^{\phi}$ to $v_{j}^{\phi}$ via an attribute node $u_{k}^{\phi}$ is:
$\color[rgb]{0,0,0}P_{i,k,j}^{AG_{\phi}}=P_{i,k}^{AG_{\phi}}\cdot P_{k,j}^{AG_{\phi}}=\bar{a}_{k,i}^{\phi}\cdot\left(1-\left|\bar{a}_{k,j}^{\phi}-\bar{a}_{k,i}^{\phi}\right|\right)$} (7)
Summing over all attribute nodes yields the total transition probability:
$\color[rgb]{0,0,0}P_{i,j}^{AG_{\phi}}=\sum_{k\in U^{\phi}}P_{i,k,j}^{AG_{\phi}}=\sum_{k\in U^{\phi}}\bar{a}_{k,i}^{\phi}\cdot\left(1-\left|\bar{a}_{k,j}^{\phi}-\bar{a}_{k,i}^{\phi}\right|\right)=\rho_{i}^{\phi}(j)$} (8)
Implication: This formalizes that AG-based random walks inherently encode type-specific semantic similarity among nodes, providing a principled mechanism to bias walks toward functionally relevant nodes.
Theorem 2 (Multi-graph Ergodicity).
Consider the multi-graph system $G_{sys}\left\{TG,AG_{o},AG_{p},AG_{l}\right\}$. Under the following assumptions:
(1) TG Connectivity: TG is strongly connected across all nodes via OD–path–segment relationships.
(2) AG internal connectivity: Each $AG_{\phi}$ is semantically connected, i.e., there exists a path between any two nodes in attribute space.
(3) Cross-graph reachability: TG ensures that any node type can reach any other node type.
(4) Mixture balance: The mixing coefficient $\alpha\in(0,1)$ ensuring probabilistic switching between TG and AG walks.
The joint random walk defines an ergodic Markov chain with a unique stationary distribution $\pi^{*}$.
Proof: The joint transition matrix for the multi-graph random walk is:
$\color[rgb]{0,0,0}P^{joint}=\alpha P^{TG}+(1-\alpha)\sum_{\phi\in\{o,p,l\}}\delta_{\phi}P^{AG_{\phi}}$} (9)
where $\delta_{\phi}=1$ if the current node is of type $\phi$ and 0 otherwise.
Irreducibility: TG connectivity ensures any node type can reach any other type via TG. AG internal connectivity guarantees that nodes of the same type are mutually reachable through attributes. Thus, any pair of nodes in $V_{sys}$ can be connected via a finite sequence of transitions, ensuring irreducibility.
Aperiodicity: Since $\alpha\in(0,1)$, the walk probabilistically switches between TG and AG, preventing fixed cycles. There exists a non-zero probability of returning to the current node (self-loop via AG), guaranteeing aperiodicity.
Positive The state space $V_{sys}$ is finite, and the chain is irreducible. Hence, every state is positive recurrent.
By the standard Markov chain theory, an irreducible, aperiodic, and positive recurrent finite chain is ergodic and possesses a unique stationary distribution $\pi^{*}$.
Implication: Regardless of the starting node, sufficiently long joint walks produce sequences whose sampling distribution converges to $\pi^{*}$, guaranteeing stability and repeatability of the embeddings learned from these sequences.
Discussion: The two theorems collectively demonstrate that TG and AGs provide complementary guidance for heterogeneous random walks. Specifically, the multiple AGs enable type-specific modeling of functional similarities, allowing transition probabilities to capture distinct functional semantics—such as OD-flow intensity, path-level relationships, and segment attributes. Meanwhile, the TG preserves multi-hop, cross-type functional dependencies, ensuring that walks capture essential functional-structural relationships among OD pairs, paths, and segments. Together, the TG and AGs form a coordinated mechanism that biases walks toward nodes that are both semantically relevant and structurally dependent. As a result, the generated heterogeneous sequences encode rich functional and structural contexts, providing a theoretical basis for learning embeddings that integrate both functional and structural dependencies.
4.2 Road segment embedding
In the heterogeneous sequences generated by HetGWalk, road segments, OD pairs, paths, and their associated attributes are interleaved in a way that often induces long-range contextual dependencies. Traditional recurrent architectures such as GRUs (gru) or LSTMs (lstm) rely on sequential inductive biases and therefore struggle to model interactions between elements that are far apart in the sequence, especially when heterogeneous entities participate in multi-hop, cross-type relationships. To effectively capture such long-range contextual dependencies, we adopt a Transformer encoder (transformer), whose fully connected self-attention mechanism directly models all pairwise interactions within each sequence without relying on positional proximity.
Before feeding a sequence into the encoder, each node is linearly projected into a type-specific embedding space, ensuring that heterogeneous entities (OD pairs, paths, segments, and attributes) are represented within a unified latent domaing. Multiple self-attention layers progressively integrate contextual information, producing context-dependent embeddings for each element. For a road segment $v_{i}^{l}$ appearing in multiple sequences, the encoder generates multiple embeddings reflecting its varying context. These embeddings are aggregated using an Attention-based Multiple Instance Learning (AMIL) mechanism (AMIL), yielding a final segment representation $h^{i}$ that adaptively weights each contextual embedding according to its importance.
$\color[rgb]{0,0,0}h^{i}=\sum_{b=1}^{\mathcal{B}}\frac{\exp\!\left\{W^{\mathrm{T}}_{\mathcal{B}1}\tanh\!\big(W_{\mathcal{B}2}h^{l}_{i,b}\big)\right\}}{\sum_{j=1}^{\mathcal{B}}\exp\!\left\{W^{\mathrm{T}}_{\mathcal{B}1}\tanh\!\big(W_{\mathcal{B}2}h^{l}_{i,j}\big)\right\}}\,h^{l}_{i,b}$} (10)
where $\mathcal{B}$ denotes the number of embeddings obtained by the transformer encoder for $v_{i}^{l}$, $W_{\mathcal{B}1}$ and $W_{\mathcal{B}2}$ are learnable parameters.
The multi-head self-attention mechanism in the Transformer captures co-occurrence patterns and functional relationships among road segments from multiple perspectives. When segments frequently co-occur within the same sequences, different attention heads specialize in recognizing distinct types of dependencies: some focus on spatial interactions induced by shared OD flows or functionally related paths, while others concentrate on attribute-driven similarities—such as traffic volume, lane count, and speed category—to identify semantic affinities. This specialized, parallel processing enables segments that serve OD pairs with comparable characteristics (e.g., similar demand levels or matching OD types) to form semantically coherent clusters in the embedding space. As a result, the final embedding for each segment effectively integrates its topological context, OD-driven long-range spatial dependencies, and diverse attribute-driven similarities, yielding a context-aware representation suitable for assessing segment criticality.
4.3 Listwise learning to rank module
We designed a listwise ranking module that predicts node importance rankings from the learned embeddings. This module consists of two stacked fully-connected layers, with a Dropout layer after the first layer to enhance generalization. It takes a list of $K$ road segment embeddings as input and outputs a list of predicted ranking scores.
To optimize this module, we adopt the well-established ListNet framework (listnet), specifically its efficient Top-One probability variant, as our training objective. This choice is motivated by its superior alignment with our goal of identifying the most critical segments. The loss function is defined as the Kullback-Leibler (KL) divergence between the ground-truth and predicted probability distributions, which are obtained by applying the SoftMax function to the ground-truth importance scores and the model’s predicted scores, respectively:
$L(Z)=D_{KL}(Y_{Z}\parallel\hat{Y}_{Z})=\sum_{v_{j}\in Z}y_{j}\log\left(\frac{y_{j}}{\hat{y}_{j}}\right)$ (11)
where $\hat{y}_{j}$ and $y_{j}$ denote the predicted ranking score and ground-truth ranking label, respectively, for the j-th road segment node in the list $Z$.
Intuitively, this loss function emphasizes the accuracy at the top of the ranking. It heavily penalizes the model if it assigns a low probability to a truly critical segment, thereby ensuring that the learning process is dominated by the most important examples.
4.4 Computational Complexity Analysis
This subsection analyzes the time complexity of HetGL2R and compares it with several representative baseline methods.
4.4.1 Complexity Analysis of HetGL2R
The computational cost of HetGL2R primarily stems from three core components:
  • 1.
    HetGWalk Sampling Module: The overall computational complexity of the joint random-walk procedure consists of a one-time preprocessing stage and the subsequent walk-generation stage. In preprocessing, Alias tables are constructed for all nodes in the Trip Graph (TG) and the Attribute-Guided Graphs (AGs), incurring a total cost of $O(m+\sum_{\phi\in\{o,p,l\}}a^{\phi})$, where $m$ is the number of TG edges and $a^{\phi}$ is the number of attribute edges in $AG_{\phi}$; The memory requirement is of the same asymptotic order. After preprocessing, each TG step requires one Alias lookup, and each AG step (entity–attribute–entity) requires two, all in $O(1)$ time. Thus, generating numnum walks of length lenlen from each of the $n$ nodes costs $O(num\cdot n\cdot len)$ in total. Overall, the joint random-walk procedure requires $O(m+\sum_{\phi\in\{o,p,l\}}a^{\phi}+num\cdot n\cdot len)$, where the first two terms correspond to preprocessing and the last term accounts for walk generation.
  • 2.
    Transformer Encoder Module: The sequences generated by HetGWalk are fed into a Transformer encoder to learn context-aware node embeddings. For sequences of length lenlen with hidden dimension $d$, the self-attention computation has complexity $O(len^{2}\cdot d)$ per sequence, and the feed-forward network contributes $O(len^{2}\cdot d)$. If numnum walks are generated for each of the $n$ nodes, the total per-epoch cost is $O(num\cdot n\cdot len^{2}\cdot d+num\cdot n\cdot len\cdot d^{2})$, with memory requirements of $O(len^{2})$ for attention maps and $O(len\cdot d)$ for embeddings per sequence.
  • 3.
    Listwise Ranking Module: This module consists of a simple feed-forward neural network with a complexity of $O(K\cdot d^{2})$, where $K$ is the ranking list length. The computational cost of this module is negligible compared to the preceding steps.
In summary, the overall complexity of HetGL2R is linear with respect to the number of sampled sequences and quadratic with respect to the sequence length. Its highly parallelizable design (especially in the walk sampling and sequence encoding stages) facilitates efficient training and inference on modern hardware (e.g., GPUs).
4.4.2 Complexity Comparison with Baseline Methods
We compare HetGL2R with two main categories of baseline methods:
  • 1.
    Message-Passing GNNs (e.g., HGT (rf48), GAT (rf50)). These models learn node representations through iterative neighborhood aggregation. Each layer incurs a computational cost of $O(|E|d)$, where $|E|$ denotes the number of edges and d the feature dimensionality. Modeling long-range dependencies requires stacking multiple layers, causing the total complexity to increase with both the graph size and the network depth. In addition, the layer-wise dependency inherent to message passing introduces sequential execution constraints, which fundamentally limit the degree of parallelism achievable within a single batch.
  • 2.
    Random Walk-based Homogeneous Methods (e.g., DeepWalk (rf39)): The sampling stage of these methods shares a similar complexity with our HetGWalk, i.e., $O(num\cdot n\cdot len)$. The key computational difference emerges in the embedding learning phase. These methods typically employ a separate Skip-gram model, which, while highly efficient per training sample due to its shallow architecture and negative sampling, operates on a corpus of generated sequences in a distinct, non-end-to-end step. In contrast, our Transformer encoder, despite having a higher per-sequence complexity of $O(len^{2}\cdot d)$, is integrated into an end-to-end framework. This design eliminates the need for a separate training pipeline and allows for the direct co-optimization of all components, although this integration introduces a higher computational cost per model update.
The computational profile of HetGL2R differs significantly from these baselines. Compared to multi-layer GNNs, our approach trades the iterative multi-hop aggregation for a more direct, sequence-based encoding of long-range contexts. Compared to pipeline-based random walk methods, it consolidates the entire learning process into a single, unified optimization, which simplifies the training workflow at the cost of increased per-pass complexity. The highly parallelizable nature of both the random walk sampling and the sequence encoding stages helps mitigate this cost in practice.
5 Experiments
HetGL2R is implemented with PyTorch. All of the experiments are conducted under a computer environment equipped with an AMD Ryzen 7 6800H processor, 16 GB of RAM, and an NVIDIA GeForce RTX 3050 Laptop GPU.
5.1 Dataset Construction
Table 2: Description of the three synthetic datasets.
   dataset    od    path    road    cars
   SY-Net110    20    1762    110    2313
   SY-Net514    90    5148    514    6334
   RD-Net3478    800    52579    3478    74174
To evaluate the proposed framework, three traffic network datasets were constructed using the SUMO (rf2) microscopic traffic simulator. Two datasets, SY-Net110 and SY-Net514, were generated from subregions of the real Shenyang urban road network, containing 110 and 514 nodes, respectively. These networks were extracted from OpenStreetMap and calibrated with realistic road capacities and speed limits. In addition, SY-Net110 and SY-Net514 incorporated real traffic flow data collected during peak hours, from which the corresponding OD flows were inferred (20 OD pairs for SY-Net110 and 90 OD pairs for SY-Net514). Furthermore, a large-scale randomly generated road network (RD-Net3478) with 3,478 nodes and 800 OD pairs was created to assess the model’s scalability to a larger network with a distinct structural topology. For all datasets, SUMO was used to simulate steady-state traffic flows, and the resulting OD–path–segment mappings were extracted to construct heterogeneous trip graphs for model training and evaluation. Table 2 summarizes the statistical properties of the three networks.
To generate the ground-truth ranking of road segment importance, we used the microscopic traffic simulation platform SUMO to simulate traffic dynamics on three networks. For each road segment, we simulated a failure scenario by reducing its maximum speed limit to 10% of its original value. The number of other segments that subsequently experienced failure (defined as a speed drop below 10% of the speed limit) was recorded over multiple time windows. The final importance score (IS (rf8)) of each segment was computed as the weighted sum of affected segments across time, capturing the overall cascading impact of its failure on the network.
$IS_{i}=\sum_{t=1}^{T}\gamma^{t}n_{t}$ (12)
where $\gamma$ is the decay coefficient (set to 0.9 in this paper), $n_{t}$ denotes the number of failed road segments in the network at time step $t$, and $T$ represents the total number of time steps.
5.2 Baselines and evaluation metrics
To evaluate the effectiveness of the proposed embedding module for ranking tasks, we compared it with the following LTR baseline methods:
  • 1.
    DeepWalk (rf39): Generates node sequences through truncated random walks and learns low-dimensional node embeddings via the Skip-gram model, preserving graph structural proximity in the latent space.
  • 2.
    GraphRNA (rf40): Efficiently integrates topological structures and node attributes through attribute-aware random walks combined with a recurrent neural architecture, enabling context-sensitive graph representation learning.
  • 3.
    TraNode2vec (rf7): Evaluates node importance by combining clustering-based topology analysis with embedding learning, effectively capturing both local and global network structures.
  • 4.
    MGL2Rank (rf8): A random walk–based attributed graph representation learning framework that employs pairwise ranking objectives to model relative node importance relationships.
  • 5.
    HGT (rf48): A heterogeneous graph transformer that models large-scale, type-rich graphs using type-dependent attention mechanisms and a scalable mini-batch sampling strategy, allowing effective learning across diverse node and edge types.
  • 6.
    Dir-GNN (rf53): Improves GNN performance on heterophilic graphs by separately aggregating incoming and outgoing edges in directed graphs while maintaining comparable accuracy on homophilic structures.
  • 7.
    GAT (rf50): Learns node representations through masked self-attention mechanisms that assign adaptive importance weights to neighboring nodes without requiring costly matrix operations.
  • 8.
    ENGINE (rf51): Generates expressive node embeddings by aggregating multi-layer hidden states from pre-trained large language models (LLMs), enabling rich semantic representation for graph-based tasks.
  • 9.
    GNNLLMEmb (rf52): Initializes node features using the last hidden layer representations from an LLM and subsequently applies a GCN to learn topology-enhanced node embeddings that combine linguistic and structural information.
  • 10.
    GATv2 (gatv2): A dynamic graph attention network that modifies the computation order of attention to make neighbor weights explicitly dependent on the query node. It achieves strictly stronger expressiveness and performance than GAT while maintaining a comparable number of parameters.
  • 11.
    PDFormer (PDFormer): A multi-view graph representation model equipped with long-range spatiotemporal attention mechanisms. It simultaneously captures local neighborhood effects and distant coupling relationships, effectively embedding congestion propagation patterns and topological constraints into the node representation space.
  • 12.
    DR-HGNN (DR-HGN): A heterogeneous road network graph model that employs dynamic relation modeling and hierarchical aggregation to jointly encode multiple node and edge types, thereby learning latent representations that integrate both structural and semantic information.
  • 13.
    SeHGNN (SeHGNN): A structure-efficient heterogeneous GNN that precomputes neighbor aggregation to reduce complexity, extends the receptive field through long meta-paths, and leverages Transformer-style semantic fusion to integrate multi-relational semantics, achieving strong representational power with efficient training.
  • 14.
    Edge-GNN (Edge-GNN): An edge-centric message-passing framework that explicitly models edge attributes and connection patterns. It aggregates information from edges to nodes, producing node representations that are more sensitive to traffic-flow propagation.
  • 15.
    LambdaLoss (LambdaLoss): A ranking-oriented loss function that directly optimizes ranking metrics by training a node scoring function to generate importance scores for ordered ranking tasks.
  • 16.
    RankFormer (Rankformer): A Transformer-based listwise learning-to-rank model that introduces listwide labels and jointly optimizes novel listwide objectives with conventional listwise losses, enabling the modeling of inter-item dependencies and improving overall ranking quality.
  • 17.
    LightGBM (lighrgbm): As a gradient boosting–based learning-to-rank baseline, LightGBM is trained with the LambdaMART objective on node attributes and topology-derived features, optimizing overall ranking quality in accordance with the evaluation metrics.
Among these methods, only MGL2Rank LambdaLoss, RankFormer and LightGBM are inherently designed for ranking. The other models are used solely to learn node embeddings, which are then fed into our proposed listwise learning-to-rank module to generate the final ranking results. According to the requirements of large language models, we convert structured or non-textual attributes into textual descriptions as inputs for ENGINE and GNNLLMEmb.
Additionally, we compared HetGL2R with the following non-learnable ranking baseline methods in terms of ranking performance:
  • 1.
    Improved PageRank (IPageRank) (rf3): Combines PageRank centrality and node attributes to rank node importance.
  • 2.
    CRRank (rf1): A heterogeneous graph-based ranking algorithm for road network nodes, considering centrality, attributes, irreplaceability, and upstream-downstream relationships between adjacent nodes.
  • 3.
    OVED-Rank (rf44): Integrating effective distance with multiple topological metrics for efficient evaluation of key spreaders’ influence in networks.
We employ the following metrics to evaluate all methods:
(1) Normalized Discounted Cumulative Gain (NDCG@K): This metric penalizes lower-ranked elements through logarithmic discounting, thereby placing greater emphasis on top-ranked items:
Table 3: Ranking Performance of Different Methods on SY-Net110.
SY-Net110 EMD NDCG@K Diff $\tau$
LTR DeepWalk 0.3146 0.7099 0.4329 0.5118
GraphRNA 0.1350 0.8145 0.3037 0.6083
TraNode2vec 0.2154 0.8011 0.3203 0.6004
MGL2Rank 0.1520 0.8970 0.2969 0.7311
HGT 0.1542 0.9287 0.2998 0.5688
Dir-GNN 0.1635 0.8847 0.3152 0.5330
GAT 0.2540 0.8757 0.3004 0.6444
ENGINE 0.2273 0.8395 0.3116 0.5984
GNNLLMEmb 0.1975 0.8617 0.3240 0.6246
GATv2 0.0889 0.9073 0.2874 0.7422
PDFormer 0.0876 0.8999 0.2772 0.7023
DR-HGNN 0.0774 0.9013 0.2685 0.6146
SeHGNN 0.0977 0.9057 0.2846 0.7592
Edge-GNN 0.1216 0.9133 0.2799 0.7805
LambdaLoss 0.0974 0.8463 0.3469 0.6413
RankFormer 0.2062 0.7656 0.4156 0.5325
LightGBM 0.1502 0.8347 0.3597 0.6557
Rank IPageRank 0.3083 0.6266 0.4658 0.4705
CRRank 0.2144 0.7797 0.3659 0.6239
OVED-Rank 0.3360 0.7578 0.3937 0.6149
HetGL2R 0.0699 0.9459 0.2483 0.7934
$\color[rgb]{0,0,0}\operatorname{NDCG@K}=\frac{\sum_{j=1}^{n}\frac{2^{n-\hat{y}_{i}}-1}{\log_{2}(j+1)}}{\sum_{j=1}^{n}\frac{2^{n-y_{i}}-1}{\log_{2}(j+1)}}$} (13)
where, $n$ denotes the list length, and $y_{i}$ represents the sorted list.
(2) Wasserstein distance (Earth Mover’s Distance,EMD): Quantifies the discrepancy between the predicted distribution and the true distribution.
(3) Ideal ranking list deviation (Diff) (rf1): Measures the relative ranking error of elements at each position in the list:
$\color[rgb]{0,0,0}\operatorname{Diff}(Y,\hat{Y})=\frac{\sum_{i=1}^{n}\left|\mathcal{P}\mathit{os}\left(v_{i}^{l},Y\right)-\mathcal{P}\mathit{os}\left(v_{i}^{l},\hat{Y}\right)\right|}{\left\lfloor n^{2}/2\right\rfloor}$} (14)
where $\mathcal{P}\mathit{os}\left(v_{i}^{l},Y\right)$ returns the order number of $v_{i}^{l}$ in the list $Y$.
(4) Kendall’s rank correlation coefficient (Kendall’s $\tau$) (tau): This metric measures the ordinal association between two rankings by comparing the number of concordant and discordant pairs:
$\color[rgb]{0,0,0}\tau\!\left(Y,\hat{Y}\right)=\frac{\operatorname{count}_{c}-\operatorname{count}_{d}}{0.5\,n\,(n-1)}$} (15)
where $\operatorname{count}_{c}$ and $\operatorname{count}_{d}$ denote the number of consistent and inconsistent pairs, respectively. The coefficient ranges from $\tau=1$ (perfect agreement) to $\tau=-1$ (perfect disagreement), with $\tau=0$ indicating no correlation. Unlike $\operatorname{NDCG@K}$, which focuses on the top-K items, Kendall’s $\tau$ evaluates the global consistency of the entire ranking.
For HetGWalk, we set the propagation factor $e$ to 2, the walk bias factor $\alpha$ to 0.6, the number of sampled sequences to 25, and the sequence length to 20. The embedding dimension of the encoding network is set to 64. The experiments are conducted on the same synthetic dataset under identical experimental settings to ensure fair comparisons.
5.3 Experimental results
For each network (SY-Net110, SY-Net514, and RD-Net3478), 70% of the nodes were randomly selected for training, and the remaining 30% were reserved for testing. Tables 3 - 5 summarize the ranking performance of all methods across multiple evaluation metrics. The results consistently show that learning-based ranking models outperform non-learning baselines across all datasets, while HetGL2R achieves the best overall performance. Specifically, HetGL2R improves $\operatorname{NDCG@K}$, Diff, and $\tau$ by at least 1.82%, 7.52%, and 1.63% on SY-Net110, by 2.98% and 4.40% on $\operatorname{NDCG@K}$ and Diff on SY-Net514 (ranking second only to HGT in $\tau$, with a marginal 0.70% gap), and by 3.39%, 4.98%, and 1.12% on SY-Net3478. These improvements confirm that the proposed embedding module effectively captures both the structural and functional dependencies of road segments, leading to robust and accurate ranking results.
Table 4: Ranking Performance of Different Methods on SY-Net514.
SY-Net514 EMD NDCG@K Diff $\tau$
LTR DeepWalk 0.2106 0.7413 0.3643 0.5005
GraphRNA 0.1103 0.8572 0.3241 0.5924
TraNode2vec 0.1822 0.8599 0.3274 0.5680
MGL2Rank 0.1357 0.9067 0.2985 0.7473
HGT 0.1309 0.8440 0.3290 0.8961
Dir-GNN 0.1381 0.8243 0.3886 0.7037
GAT 0.1371 0.8687 0.3001 0.6905
ENGINE 0.1723 0.8623 0.3109 0.5934
GNNLLMEmb 0.1840 0.8759 0.2988 0.6012
GATv2 0.2195 0.8754 0.3070 0.6901
PDFormer 0.1956 0.9189 0.2659 0.8226
DR-HGNN 0.1964 0.9118 0.2693 0.7893
SeHGNN 0.1961 0.9007 0.2946 0.7491
Edge-GNN 0.1952 0.9163 0.2592 0.8007
LambdaLoss 0.1799 0.8840 0.3201 0.6915
RankFormer 0.1942 0.8685 0.3624 0.6178
LightGBM 0.2240 0.8890 0.3094 0.7390
Rank IPageRank 0.2443 0.6573 0.4446 0.4857
CRRank 0.1781 0.7464 0.4089 0.5275
OVED-Rank 0.2833 0.7809 0.3967 0.5890
HetGL2R 0.1318 0.9471 0.2478 0.8898
Table 5: Ranking Performance of Different Methods on RD-Net3478.
RD-Net3478 EMD NDCG@K Diff $\tau$
LTR DeepWalk 0.4638 0.6525 0.4850 0.3114
GraphRNA 0.3209 0.7545 0.3843 0.5229
TraNode2vec 0.5175 0.7428 0.3957 0.4535
MGL2Rank 0.3938 0.8066 0.3492 0.6629
HGT 0.4076 0.7630 0.3781 0.5679
Dir-GNN 0.4114 0.7405 0.4042 0.4567
GAT 0.4167 0.7394 0.4176 0.3974
ENGINE 0.4180 0.7610 0.3923 0.5348
GNNLLMEmb 0.4197 0.7751 0.3987 0.3854
GATv2 0.4043 0.7904 0.3742 0.6524
PDFormer 0.3265 0.8495 0.3441 0.6990
DR-HGNN 0.3504 0.8196 0.3679 0.7361
SeHGNN 0.3517 0.8214 0.3796 0.6817
Edge-GNN 0.3554 0.8160 0.3773 0.7829
LambdaLoss 0.3476 0.7858 0.3842 0.5836
RankFormer 0.3672 0.7740 0.4170 0.5680
LightGBM 0.4686 0.7552 0.3769 0.5323
Rank IPageRank 0.4431 0.6029 0.5035 0.2254
CRRank 0.3802 0.7591 0.4074 0.5006
OVED-Rank 0.3923 0.6948 0.4510 0.4050
HetGL2R 0.3208 0.8502 0.3318 0.7918
Figure 4: The Performance of Each Method on SY-Net110.
Figure 5: The Performance of Each Method on SY-Net514.
Figure 6: The Performance of Each Method on RD-Net3478.
Table 6: Predicted Rankings of the Top 10 Nodes in the Ground Truth of SY-Net110.
DeepWalk GraphRNA TraNode2vec MGL2Rank HGT Dir-GNN GAT ENGINE GNNLLMEmb GATv2 PDFormer
2 3 1 1 1 1 1 1 6 1 1
3 7 9 2 2 2 6 5 8 2 2
5 5 4 3 7 3 3 2 3 3 3
4 2 3 4 5 4 4 3 5 4 4
1 4 2 5 6 16 2 7 7 5 5
8 8 8 6 3 8 10 10 4 6 6
11 13 19 9 8 7 16 12 2 11 9
7 6 6 7 10 5 8 4 27 7 7
19 10 31 8 4 9 25 32 19 8 8
6 1 10 11 20 6 9 8 28 10 10
DR-HGNN SeHGNN Edge-GNN LambdaLoss RankFormer LightGBM IPageRank CRRank OVED-Rank HetGL2R Ground-Truth
1 1 1 1 1 2 2 1 1 1 1
3 2 2 2 2 1 4 2 2 2 2
2 3 3 6 3 8 1 3 3 3 3
4 4 4 3 4 5 5 4 4 4 4
5 10 5 4 5 7 3 5 7 5 5
6 13 6 12 6 6 7 11 11 7 6
12 7 9 19 13 4 21 7 8 6 7
7 9 7 5 7 19 9 10 10 8 8
9 8 8 9 11 20 16 9 9 10 9
11 12 10 7 12 12 17 12 14 9 10
To provide an intuitive comparison of algorithm performance, Figures 4 - 6 illustrate the correspondence between predicted rankings and ground truth importance scores (IS) across the SY-Net110, SY-Net514, and RD-Net3478 datasets. Overall, as predicted rankings decrease, the ground truth IS values for all methods exhibit the expected monotonic downward trend. In the two simulated real-world road networks, HetGL2R demonstrates the smoothest and most stable decline, with minimal scatter, closely matching the ground truth and showing excellent ranking consistency. MGL2Rank, SeHGNN, PDFormer, and Edge-GNN also display reasonable trends, though with slightly larger overall fluctuations. In the randomly generated RD-Net3478 network, HetGL2R maintains robust performance with tightly clustered and orderly scatter points. In contrast, DeepWalk, GAT, GNNLLMEmb, and IPageRank show more dispersed scatter and greater deviation from the true rankings. Structure-driven models such as HGT, GAT, and Dir-GNN perform adequately on the simulated real networks but experience significant performance drops on the randomized network. We attribute this performance difference to the weaker coupling between network topology and OD flows in the synthetic random network. In real-world urban road networks, topology and OD flows are typically strongly coupled: central or highly connected segments are more likely to carry significant OD flows, allowing structural features to serve as reliable proxies for importance. In the RD-Net3478 network, however, this coupling is substantially weaker, making it difficult for topology-dependent models to accurately infer criticality. Furthermore, as a general-purpose heterogeneous graph learning model, HGT is limited by its primary reliance on node and edge types, making it difficult to fully leverage rich edge attributes in path sequences (such as node order), which are crucial for capturing traffic flow directionality and influence scope. HetGL2R serializes the graph structure via heterogeneous random walks and explicitly captures sequential dependencies using the Transformer’s full-sequence modeling capabilities, outperforming message-passing-based heterogeneous GNNs in representing road network structures. Notably, PDFormer achieves high-level performance across all three networks. As a state-of-the-art method designed for traffic networks, it employs an attention mechanism guided by semantic masks to focus on geographically distant but functionally similar node pairs, capturing long-range dependencies, while enhanced key computations simulate propagation delay. Its approach aligns conceptually with HetGL2R in capturing functional and structural dependencies, but differs technically. HetGL2R explicitly models OD flows, paths, and inter-segment relationships to learn more precise long-range dependencies, resulting in more robust performance in experiments. Graph-based large language models have yet to surpass specialized graph learning methods, primarily because they are designed for general graph structures and do not effectively incorporate OD flows, which fundamentally determine long-range spatial dependencies, thereby hindering their ability to accurately capture such dependencies.
Table 7: Predicted Rankings of the Top 10 Nodes in the Ground Truth of SY-Net514.
DeepWalk GraphRNA TraNode2vec MGL2Rank HGT Dir-GNN GAT ENGINE GNNLLMEmb GATv2 PDFormer
22 1 1 1 1 1 1 69 1 1 1
25 4 5 5 3 4 5 12 2 3 5
1 3 2 2 2 3 2 1 7 2 2
8 2 4 4 4 2 4 6 69 4 4
2 12 3 3 5 5 3 2 33 5 3
7 13 6 6 6 6 6 5 4 6 6
55 7 9 12 7 7 8 44 10 7 9
13 10 7 7 8 9 7 8 9 8 7
45 9 12 13 9 11 10 30 14 11 10
46 17 14 14 10 13 11 31 3 10 11
DR-HGNN SeHGNN Edge-GNN LambdaLoss RankFormer LightGBM IPageRank CRRank OVED-Rank HetGL2R Ground-Truth
1 1 2 2 14 1 21 3 1 1 1
3 4 5 6 24 5 2 1 2 3 2
4 7 1 1 43 3 51 7 7 2 3
20 8 4 4 4 2 53 4 9 4 4
2 6 3 3 51 4 20 2 4 5 5
5 2 6 5 17 7 7 5 5 6 6
6 5 11 14 10 8 60 24 15 7 7
7 3 7 7 2 6 44 9 11 8 8
13 10 13 12 7 12 10 17 10 9 9
10 15 14 13 8 11 70 21 33 10 10
Table 8: Predicted Rankings of the Top 10 Nodes in the Ground Truth of RD-Net3478.
DeepWalk GraphRNA TraNode2vec MGL2Rank HGT Dir-GNN GAT ENGINE GNNLLMEmb GATv2 PDFormer
237 3 466 1 1 96 5 184 376 12 1
443 19 41 4 6 50 34 16 1 2 4
610 4 107 2 3 2 4 50 4 6 2
3 1 217 3 2 31 3 101 7 15 3
1 86 665 19 37 8 135 312 5 54 18
612 56 18 16 22 23 81 8 35 3 15
599 11 169 5 10 117 15 88 27 19 5
409 20 46 11 15 26 31 28 15 7 10
405 15 22 9 12 82 18 12 242 4 8
187 47 49 18 30 87 62 29 8 11 19
DR-HGNN SeHGNN Edge-GNN LambdaLoss RankFormer LightGBM IPageRank CRRank OVED-Rank HetGL2R Ground-Truth
14 3 6 18 1 3 385 4 15 2 1
2 1 1 11 2 2 445 1 4 1 2
6 7 4 1 4 1 568 12 6 3 3
15 20 9 9 7 5 3 2 2 5 4
52 5 27 3 9 4 2 5 14 9 5
3 26 3 8 5 6 310 19 11 4 6
19 66 15 46 13 14 458 16 3 10 7
7 21 7 13 12 7 494 21 20 7 8
4 191 5 37 10 12 486 34 21 6 9
10 15 10 43 17 19 247 15 36 11 10
In the two simulated real-world road networks, HetGL2R demonstrates the smoothest and most stable decline, with minimal scatter, closely matching the ground truth and showing excellent ranking consistency. MGL2Rank, SeHGNN, PDFormer, and Edge-GNN also display reasonable trends, though with slightly larger overall fluctuations. In the randomly generated RD-Net3478 network, HetGL2R maintains robust performance with tightly clustered and orderly scatter points. In contrast, DeepWalk, GAT, GNNLLMEmb, and IPageRank show more dispersed scatter and greater deviation from the true rankings. Structure-driven models such as HGT, GAT, and Dir-GNN perform adequately on the simulated real networks but experience significant performance drops on the randomized network. We attribute this difference to the weakened correlation between network topology and OD flows in the random network. In real urban road networks, topology and OD flows are typically coupled: central or highly connected segments are more likely to carry major OD flows, allowing structural features to provide strong cues for node importance assessment. In RD-Net3478, however, this topology–OD coupling is weakened, making it difficult for models that rely solely on topology to reliably infer node importance.
Tables 6 to 8 present the predicted rankings of the ground-truth top-10 nodes across the three networks. Comprehensive analysis demonstrates that HetGL2R achieves the highest ranking consistency and PDFormer delivers competitive performance. Both models exhibit exceptional generalization capability and robustness across different network configurations.
Figure 7: Case study on segment ranking in SY-Net514 simulated road network.
Table 9: Runtime and Memory Usage of HetGL2R and Baseline Methods on Three Networks.
Method Time(s)
SY-Net110 SY-Net514 RD-Net3478
DeepWalk 7.12 41.93 299.71
GraphRNA 17.83 149.82 326.18
TraNode2vec 8.15 56.12 378.29
MGL2Rank 13.83 123.63 302.25
HGT 17.54 104.89 190.25
Dir-GNN 16.77 108.86 211.84
GAT 12.03 124.19 311.47
ENGINE 15.18 217.56 882.36
GNNLLMEmb 28.40 300.18 680.40
GATv2 15.42 147.18 412.34
PDFormer 125.21 321.43 1090.57
DR-HGNN 152.33 473.79 1544.77
SeHGNN 31.61 115.87 341.36
Edge-GNN 128.89 381.57 1130.66
LambdaLoss 9.31 13.09 122.63
RankFormer 11.28 16.96 163.25
LightGBM 2.56 4.13 13.27
HetGL2R 16.27 126.93 314.92
Figure 7 illustrates two representative road segments during ranking evaluation on the SY-Net514 network. In the rankings generated by HetGL2R, segment #470 has a relatively low traffic volume (215 vehicles/hour) but is ranked 15th, higher than segment #336, which carries a larger traffic volume (588 vehicles/hour, ranked 22th). Further analysis shows that segment #470 is located near the destination of a certain OD pair, with upstream nodes carrying substantial traffic flow. A failure at segment #470 could block upstream traffic, causing congestion in the upstream region and potentially impacting the broader network performance. This example demonstrates that HetGL2R’s ranking considers not only node attributes or local topology but also the structural dependencies along upstream and downstream paths.
5.4 Computational Efficiency and Scalability Analysis
To evaluate the scalability of HetGL2R with respect to network size, we analyzed its per-epoch running time on three networks of different scales (Table 9). Our key findings are as follows: (i) Compared to computationally lightweight methods (e.g., DeepWalk, TraNode2vec), specialized ranking models (e.g., LambdaLoss, RankFormer), and the lightweight ensemble model LightGBM, HetGL2R requires significantly longer training time due to its more complex architecture; (ii) However, when compared against other graph neural network models (e.g., PDFormer, DR-HGNN, Edge-GNN), HetGL2R demonstrates comparable efficiency. This comparable efficiency is associated with its random walk sampling strategy. This strategy confines the computational complexity to a linear relationship with the walk length and sampling count, thereby avoiding the rapid increase in computational cost caused by the quadratic or higher-order polynomial complexity with respect to the number of nodes, as seen in methods employing global graph attention (e.g., PDFormer) or deep hierarchical message passing (e.g., DR-HGNN). Regarding scalability, the primary bottlenecks for HetGL2R lie in the Transformer encoder’s processing of long sequences and the storage I/O required for full-graph walk sampling. Consequently, its scalability on very large graphs is still expected to surpass that of most GNNs, yet may not match that of lightweight models requiring no graph structure. In summary, through its core design of trading sampling for computation, HetGL2R achieves an effective balance between the complex requirements of heterogeneous graph ranking and computational feasibility.
Table 10: Ablation Study on Ranking Performance under four Views.
module SY-Net110     SY-Net514     RD-Net3478
EMD NDCG@K Diff $\tau$ EMD NDCG@K Diff $\tau$ EMD NDCG@K Diff $\tau$
w/o AG 0.2198 0.9071 0.2811 0.7284 0.1224 0.8760 0.3053 0.7919 0.3275 0.7774 0.3806 0.7117
w/o transformer(GRU) 0.1564 0.7882 0.3813 0.6103 0.1698 0.7714 0.3761 0.7866 0.4048 0.6569 0.4579 0.5906
w/o transformer(LSTM) 0.2700 0.7646 0.4021 0.6829 0.2695 0.7693 0.3992 0.7703 0.3746 0.7393 0.3930 0.6634
W/o AMIL(Mean) 0.0637 0.9116 0.2536 0.7335 0.1221 0.9380 0.2404 0.8638 0.3882 0.8236 0.3402 0.7302
W/o AMIL(DeepSets) 0.0788 0.9215 0.2533 0.7432 0.1474 0.9252 0.2645 0.8591 0.3726 0.8344 0.3461 0.7574
W/o AMIL(SetTrans) 0.0705 0.9246 0.2834 0.7724 0.1421 0.9458 0.2648 0.8741 0.3940 0.8468 0.3612 0.7759
W/o KL(ListNet) 0.0761 0.9190 0.2814 0.7485 0.2607 0.9054 0.2744 0.8478 0.3512 0.8034 0.3526 0.7462
W/o KL(ListMLE) 0.0939 0.9108 0.2830 0.7341 0.2340 0.8483 0.3655 0.7208 0.3670 0.7467 0.4018 0.7028
HetGL2R 0.0699 0.9459 0.2483 0.7934 0.1318 0.9471 0.2478 0.8898 0.3208 0.8502 0.3318 0.7918
5.5 Ablation Study
We conducted ablation studies on HetGWalk, the Transformer encoder, the AMIL aggregation function, and the KL-divergence loss, with results shown in Table 10.
Ablation Study on HetGWalk. We evaluated the impact of joint random walks on the Trip Graph (TG) and Attribute-guided Graph (AG) by comparing the full model with a variant excluding AG (w/o AG). The w/o AG variant consistently exhibits substantial performance drops across all networks and metrics, indicating that joint random walks are essential for achieving high ranking accuracy and robustness. This decline arises from TG’s limitations: it captures functional structural dependencies but cannot model functional similarity, and relying solely on topology biases node sampling toward high-centrality nodes while underexploring latent functional clusters. AG mitigates these issues by enabling transitions between semantically similar nodes, enriching contextual diversity. Consequently, the joint TG–AG random walk effectively integrates structural and functional perspectives, generating context-rich sequences that support robust and accurate segment representations.
Ablation on Transformer Encoder. We replaced the Transformer with GRU (gru) and LSTM (lstm) (denoted as w/o transformer (GRU) and w/o transformer (LSTM), respectively). Experimental results show that both RNN variants performed significantly worse than the Transformer architecture across all datasets and metrics. The performance degradation stems from the limitations of RNNs’ sequential modeling mechanism in capturing long-range dependencies, while Transformer’s global self-attention can directly model high-order dependencies between any positions in the sequence, making it more suitable for processing complex sequences generated by heterogeneous graph random walks.
Ablation Study on AMIL aggregation. We performed ablation experiments on the context embedding aggregation strategy by replacing AMIL with Mean, DeepSets (deepsets), and SetTrans (SetTrans) aggregation, denoted as w/o AMIL (Mean), w/o AMIL (DeepSets), and w/o AMIL (SetTrans), respectively. AMIL consistently achieves the best trade-off between performance and efficiency, attaining the highest NDCG@K scores across all datasets and outperforming alternatives on Diff and EMD metrics. The results indicate that Mean fails to differentiate critical contexts, DeepSets lacks inter-instance interaction, and SetTrans overemphasizes fine-grained variations, whereas AMIL’s lightweight attention mechanism effectively captures key contextual information while mitigating noise. These findings validate the effectiveness and superiority of AMIL as the aggregation strategy in our framework.
Ablation study on the KL divergence loss. we replaced the KL divergence with the cross-entropy loss used in ListNet (listnet) and the likelihood loss used in ListMLE (mle), denoted as w/o KL (ListNet) and w/o KL (ListMLE), respectively. Results show that KL divergence outperforms the other losses across all metrics. This reflects the conceptual differences among listwise losses: ListMLE maximizes the likelihood of the true ranking, but its stepwise generation assumption can accumulate errors in noisy traffic network scenarios; ListNet optimizes only the Top-one probability and fails to fully exploit the ranking information of the rest of the list. In contrast, KL divergence performs a soft alignment between the true and predicted importance distributions, with the true distribution (generated by applying Softmax to the true scores) naturally assigning higher weights to the top-ranked nodes, thereby focusing the training on accurate top-ranking predictions. Moreover, KL divergence does not rely on stepwise sequence generation, enabling robust handling of ranking uncertainty and fully utilizing supervision signals from all positions in the list. Based on these properties, KL divergence achieves the best performance in our task.
Figure 8: Parameter sensitivity analysis.
Comprehensive ablation experiments prove that each core component in the HetGL2R framework contributes significantly to the final performance. The organic combination of these components ensures the model’s superior performance in critical road segment identification tasks.
5.6 Sensitivity analysis
We conducted a sensitivity analysis on key hyperparameters, including the walk bias factor $\alpha$, the sampling sequence length lenlen, the list length $K$, and the depth-first exploration probability $\beta$, with results summarized in Figure 8. When $\alpha$ is set to 0.6, the model achieves the highest Kendall $\tau$, indicating an effective balance between node topology and attribute information. The optimal sampling length lenlen is 20, as shorter sequences fail to capture sufficient contextual information from related nodes, while longer sequences may introduce noise that degrades performance. Similarly, a list length of $K=5$ allows HetGL2R to best capture intra-list ranking dependencies among ranked elements; shorter lists lack internal coherence, whereas longer lists increase task complexity and dilute the ranking signal. Finally, setting $\beta$ to 0.8 yields the highest Kendall $\tau$, showing that a moderately strong preference for depth-first exploration balances exploitation of long-range structural dependencies with control of noise. Values of $\beta$ below 0.6 underutilize multi-hop context and reduce ranking concordance, whereas $\beta=1.0$ effectively switches to pure breadth-first exploration, weakening deep relational signals and sharply decreasing performance. Overall, these results highlight the sensitivity of HetGL2R to these hyperparameters and provide guidance for configuring them to achieve optimal ranking accuracy.
6 Conclusion and Future Work
This paper proposes HetGL2R, a heterogeneous graph learning framework for traffic network ranking. The framework employs heterogeneous random walks and Transformer architecture to jointly model the functional relationships among OD flows, paths, and road segments along with their attribute similarities. By integrating functional and structural information, it effectively captures long-range spatial dependencies, generating context-aware road segment embeddings. These embeddings are subsequently processed by a listwise ranking module to produce the final importance ranking. Extensive evaluations on three simulated networks validate the framework’s superior accuracy for ranking road networks.
This study assumes consistent OD–path–segment mappings derived from accurate OD demand and route choice data. In practice, uncertainties in OD estimation and potential errors in path enumeration may affect the fidelity of the constructed heterogeneous trip graph. Moreover, the current evaluation relies on SUMO-based simulated networks, which provide controlled yet idealized settings. Future work will focus on enhancing robustness to noisy or incomplete OD data, integrating uncertainty modeling, and validating the framework on large-scale real-world traffic networks.
Statement
During the preparation of this work the author(s) used DeepSeek and GPT4 in order to enhance the linguistic fluency of the manuscript. After using this tool/service, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the publication.